OpenStudy (k12girlohva):

MEDAL AND FAN Ethan needs to order T-shirts for all the members. Each T-shirt costs $15 and the shipping fee for the order is $9.99. (a) Write an equation showing the total cost, y, for an order of x T-shirts. (b) Ethan has $500 allocated in the club’s budget for T-shirt orders. How many T-shirts can Ethan order with this amount? Show your work and explain your answer. (c) There are 45 members in the club. How much money would Ethan need in order to purchase each member 1 T-shirt? Show your work

OpenStudy (k12girlohva):

MEDAL AND FAN!!

OpenStudy (k12girlohva):

@Mehek14

OpenStudy (k12girlohva):

Any1

OpenStudy (anonymous):

So I know the answer is C

OpenStudy (k12girlohva):

@Wifi-Jock could u explain?

OpenStudy (anonymous):

Yes hold on, I'm trying to give you my screenshot of my work.

OpenStudy (k12girlohva):

Ok thanks

OpenStudy (qwertty123):

HEY, 15t + 9.99 <= 500 -- subtract 9.99 from both sides 15t <= 500 - 9.99 15t <= 490.01 -- divide both sides by 15 t <= 490.01/15 So what is that?

OpenStudy (k12girlohva):

THANKS

OpenStudy (qwertty123):

You Welcome! :D Do you know what the answer is?

jhonyy9 (jhonyy9):

a) you need writing an equation for a total cost y for an order of x shirts try it for 1 shirt - so you know that 1 shirt will cost $15 + $9,99 - yes ? so how you think that will can you rewrite it for x numbers of shirt ?

jhonyy9 (jhonyy9):

@Qwertty123 every exercises start to solve them by a. - or not ?

OpenStudy (qwertty123):

Yes sorry! :D Or Step 1!

OpenStudy (qwertty123):

@jhonyy9

jhonyy9 (jhonyy9):

@Qwertty123 come please this is your work - you have got for here a medal - so please try finish your help here - ok ? thank you very much

OpenStudy (qwertty123):

He got the answer... I am so confused?

jhonyy9 (jhonyy9):

for a. ,b. and c. - i have thought your help

OpenStudy (qwertty123):

OH!!! OMW I just saw that!

OpenStudy (qwertty123):

HEY, 15t + 9.99 <= 500 -- subtract 9.99 from both sides 15t <= 500 - 9.99 15t <= 490.01 -- divide both sides by 15 t <= 490.01/15 So what is that? Then.. check.. 15(32) + 9.99 <= 500 480 + 9.99 <=?...

OpenStudy (k12girlohva):

489.99

OpenStudy (qwertty123):

YEAH! Okay... check.. 15(32) + 9.99 <= 500 480 + 9.99 <= 500 489.99 <= 500 (correct) Ethan can order 32 or less t-shirts

OpenStudy (k12girlohva):

Thanks:)

OpenStudy (qwertty123):

You Welcome! :D